OpenStudy (anonymous):

Solve for x: x^2=x+4

OpenStudy (anonymous):

bring (x+4) to the other side and either factor or use the quadratic equation.

OpenStudy (lgbasallote):

put everything on one side first x^2 - x - 4 = 0 now use the quadratic formula \[\large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[\large x = \frac{-(-1) \pm \sqrt{(-1)^2 - 4(1)(-4)}}{2(1)}\] got it?

OpenStudy (anonymous):

Is there a way of knowing when you have to use the quadratic equation and when you can just factor?

OpenStudy (lgbasallote):

well i for one initially thought it can be factored but i couldnt...my point is i base my judgment on trial and error...

OpenStudy (anonymous):

Theres not really a method, you have to eyeball it. If you can't think of numbers off the top of your head or it is not obvious what the answer should be, then most likely it will be using the Quadratic Equation

OpenStudy (lgbasallote):

but if you think abt it the factors of 4 are: 4 x 1 and 2 x 2 neither factor can produce -1 when added/subtracted

OpenStudy (anonymous):

Thanks :)